OpenStudy (anonymous):

How to find x? 2^2 * 2^x * 2^3 = 8

OpenStudy (studygurl14):

\(\large 2^22^x2^3=8\) First step, find the value of \(2^2\) and \(2^3\)

OpenStudy (confxsedaf):

add the powers: 2^2+x+3=8 you get 2^x+5=8 Then convert both sides to the same base 2^x+5=2^3 Then subtract the base of 2 on both sides x+5=3 Subtract 5 both sides x=3-5 Last simplify 3-5 to -2 x=-2 hope i helped!

OpenStudy (anonymous):

Ohh thank u, it isn't that hard xD

OpenStudy (confxsedaf):

welcome. nope good luck